OpenStudy (anonymous):

What value of m solves the equation? 2^m = 1/8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@pooja195

OpenStudy (mathstudent55):

Have you learned negative exponents?

OpenStudy (anonymous):

doing it now .-. most annoying thing ever.

OpenStudy (mathstudent55):

\(\large a^{-n} = \dfrac{1}{a^n} \)

OpenStudy (mathstudent55):

First, rewrite the fraction 1/8 as 1 over a power of 2. 8 is 2 to what power?

OpenStudy (mathstudent55):

\(\large 2^? = 8\)

OpenStudy (anonymous):

I dont know. ;-;

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

\[2 \times 2 = ?\]

OpenStudy (mathstudent55):

\(8 = 2 \times 2 \times 2\) Ok?

OpenStudy (mathstudent55):

How do you write \(2 \times 2 \times 2\) as 2 raised to an exponent?

OpenStudy (mathstudent55):